On March 1, 2019, Annapolis Company has a beginning Work in Process inventory of zero. All materials are added into production at the beginning of its production. There is only one production WIP inventory. During the month 33,000 units were started. At the end of the month all started units were 80% complete with respect to conversion. Direct Materials placed into production had a total cost of $500,000 and the total conversion cost for the month was $473,000. Annapolis uses the weighted-average process costing method. Use this information to determine the cost per equivalent unit of direct material for the month of March. (Round answer to the nearest cent.)
Answer #1

Equivalent units for direct materials = 33,000*100%= 33,000

Equivalent units for conversion= 33,000*80%= 26,400

Equivalent cost per unit for direct materials= 500,000/33,000= 15.15

Equivalent cost per unit for conversion = 473,000/26400= 17.92

Total cost per equivalent unit = 15.15+ 17.92 = $33.07